Just before a two stage rocket separates, the velocity of the combined sections is 2800 m/s. Section A has a mass of 850 kg and section B has a mass of 4.1 tonne. Both sections travel on in the same direction as before, but section A, which accommodates the astronauts, is moving 305 m/s faster than section B.

Calculate the final velocity of section A
Answer #1

the final velocity of section A will be

(v_{f})_{A}=u.ln(\frac{M_{i}}{M_{f}})+v_{i}

where u is the relative velocity of section A w.r.t section B

M_{i} is initial mass of rocket, (i.e. combined stages A+B)

M_{f} is the final mass of rocket, (i.e. stage A)

v_{i}  is the initial velocity of rocket just before separation of stages

so (v_{f})_{A}=305.ln(\frac{(4100+850)}{850})+2800

(v_{f})_{A}=305.ln(\frac{4950}{850})+2800

(v_{f})_{A}=3337.38m/s